# Detection of a Type-C QPO during the soft-to-hard transition in Swift J1727.8 –1613

### v1

@@ 5 Summary
To summarize, we identified a Type-C QPO in the soft–to–hard transition of the BH LMXB Swift J1727.8–1613, significant in both the PS and CS. As the source transitions from the soft to the hard state, the QPO frequency shifts to lower frequencies, the broadband fractional rms amplitude in the 0.3–2 keV energy band increases, and the rms covariance of the QPO decreases. At the QPO frequency, the phase lags increase and the coherence function drops, probably due to variations in the corona geometry. Moreover, XMM-Newton’s unprecedented detection of these features in the elusive decaying phase of the outburst support a physical origin of the low-frequency excess observed in the lags and coherence. Follow-up modelling of this source may help understand the relative strength of the Type-C QPO in the Imaginary component of the CS.
@@ Acknowledgements.
Research is based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. M.B. acknowledges the support from GAUK project No. 102323. J.S. thanks GACR project 21-06825X for the support. NCS acknowledges support from the Science and Technology Facilities Council (STFC) grant ST/X001121/1. FG acknowledges support from PIBAA 1275 and PIP 0113 (CONICET). FG was also supported by grant PID2022-136828NBC42 funded by the Spanish MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe. MM acknowledges the research programme Athena with project number 184.034.002, which is (partly) financed by the Dutch Research Council (NWO). We thank Santiago del Palacio for valuable discussion and comments.
@@
Appendix A
This data set is part of a multiwavelength observational campaign with Epochs defined by the optical observations from the Very Large Telescope (VLT), performed with the X-Shooter spectrograph (Vernet et al., 2011), see Castro Segura et al. in preparation.
@@
Figure 5: Hardness intensity diagram (HID) of Swift J1727.8–1613 from the 2023 outburst. The photon fluxes are derived from the MAXI/GSC instrument with a binsize of 1 day from the ”on-demand process” archives. The hardness ratio is defined as the photon flux ratio in the 4–10 keV over the photon flux in the 2–4 keV energy band. The pink squares indicate the beginning of the radio flaring period of the source and when the radio flux remains constant around 0.8 mJy (see radio monitoring of the source from Hughes et al. submitted), and the purple crosses indicate the position of the three radio flares.
Fig. 5 shows the hardness intensity diagram (HID) of Swift J1727.8–1613 since the start of the 2023 outburst. The light curves were obtained from the Monitor of All-sky X-ray Image’s (MAXI; Matsuoka et al., 2009) Gas Slit Camera (GSC; 2–30 keV; Mihara et al., 2002; Sugizaki et al., 2011; Mihara et al., 2011), in the 2–10 keV energy band, and the MAXI ”on-demand process” archives [^2] http://maxi.riken.jp/mxondem. The position of the source in the HID during the XMM EPIC-PN observations are shown in black diamonds. The hardness ratio (HR) is defined as the ratio of the photon flux in the 4–10 keV energy band to that in the 2–4 keV energy band. While the pink squares indicate the beginning of the radio flaring period of the source and when the radio flux remains constant around 0.8 mJy (see radio monitoring of the source from Hughes et al. submitted), the purple crosses highlight the position of the three radio flares within the flaring period.

### v2

@@ 5 Summary
    This analysis revealed a type-C QPO in the soft–to–hard transition of the LMXB Swift J1727.8–1613. This represents the first ever detection of a type-C QPO in the soft–to–hard transition of a BH XB using XMM–Newton data (it was previously only observed with other telescopes, e.g., NICER or RXTE).
    The QPO is significantly detected in the imaginary part of the CS and the PS, with a central frequency at $\approx$2.6 Hz, shifting to $\approx$2 Hz as the source spectrum hardens. This indicates geometrical changes in the corona. The broadband fractional rms amplitude in the 0.3–2 keV energy band increases simultaneously and the rms covariance of the QPO decreases, which suggests fluctuations in the accretion disk (probably from the mass-accretion rate) and modulated by the Comptonization. Considering the total rms of about 17% and 30% between the last two observations and the brightness of the source, we classified the QPO as a type-C QPO.
    The unique detection of the coherence drop and the phase-lag excess at the QPO frequency in J1727 suggests a weaker interaction between the large corona and the accretion disk at the QPO frequency as the transition evolves from the soft to the hard state. The first detection of the coherence drop in the soft–to–hard transition with a telescope other than NICER implies that it has a physical origin and is not an instrumental artefact.
Follow-up modeling of this source may help us to understand the relative strength of the type-C QPO in the imaginary component of the CS.
@@ Acknowledgements.
Research is based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. M.B. acknowledges the support from GAUK project No. 102323. J.S. thanks GACR project 21-06825X for the support. NCS acknowledges support from the Science and Technology Facilities Council (STFC) grant ST/X001121/1. FG acknowledges support from PIBAA 1275 and PIP 0113 (CONICET). FG was also supported by grant PID2022-136828NBC42 funded by the Spanish MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe. MM acknowledges the research programme Athena with project number 184.034.002, which is (partly) financed by the Dutch Research Council (NWO). FMV acknowledges support from the European Union’s Horizon Europe research and innovation programme with the Marie Skłodowska-Curie grant agreement No. 101149685. We thank Santiago del Palacio for valuable discussion and comments.
@@
Appendix A Observation log and light curve
@@
This dataset is part of a multiwavelength observational campaign with epochs defined by the optical observations from the Very Large Telescope (VLT), performed with the X-shooter spectrograph (Vernet et al. 2011); see Castro Segura et al. in preparation.
@@
Appendix B Hardness intensity diagram
Figure 5: Hardness intensity diagram (HID) of Swift J1727.8–1613 from the 2023 outburst. The photon fluxes are derived from the MAXI/GSC instrument with a bin size of 1 day from the ”on-demand process” archives. The hardness ratio is defined as the photon flux ratio in the 4–10 keV over the photon flux in the 2–4 keV energy band.
Figure 5 shows the hardness intensity diagram (HID) of Swift J1727.8–1613 since the start of the 2023 outburst. The light curves were obtained from the Monitor of All-sky X-ray Image’s (MAXI; Matsuoka et al. 2009) Gas Slit Camera (GSC; 2–30 keV; Mihara et al. 2002; Sugizaki et al. 2011; Mihara et al. 2011), in the 2–10 keV energy band, and the MAXI ”on-demand process” archives [^3] http://maxi.riken.jp/mxondem. The position of the source in the HID during the XMM EPIC-PN observations are shown in black diamonds. The hardness ratio (HR) is defined as the ratio of the photon flux in the 4–10 keV energy band to that in the 2–4 keV energy band.